pipe A fills a certain swimming pool in 5 hours. Pipe B fills the swimming pool in 3 hours and Pipe C drains it in 10 hours. Pipe A and Pipe B were turned on to fill the swimming pool but pipe C was accidentally left open. How long will it take for Pipe A and Pipe B to fill the swimming pool while Pipe C was left open?
Answer by macston(5194)

Pipe A fills 1pool/5hrs=1/5 pool/hr
Pipe B fills 1pool/3hrs=1/3 pool/hr
Pipe C fills -1 pool/10 hrs=-1/10 pool/hr
With all three open:
Pipe A + Pipe B + Pipe C=1/5+1/3-1/10=6/30+10/30-3/30=13/30 pool per hour
The pool fills at a rate of 13/30 of the pool per hour with all pipes open
A t this rate 1 pool takes 1pool/(13/30)pool/hr=2.3 hrs
ANSWER it takes 2.3 hours to fill the pool with all pipes open